OpenStudy (anonymous):

Kyra observed that 85% of the students at a library liked reading mystery stories. If 30 students at the library did not like reading mystery stories, the number of students at the library who liked mystery stories was _____. (only put numeric values, no other symbols)

OpenStudy (azureilai):

\[\frac{ 30 }{ x }=\frac{ 15 }{ 100 }\] Solve for x so 15x=(30)(100) then you subtract x from the solution you got from above to find the number of students that liked reading mysteries.

OpenStudy (anonymous):

@Azureilai thanks love